OpenStudy (anonymous):

Given that, \[\int\limits_{-1}^{4} f(x) dx = 3 \] find: \[\int\limits_{-1}^{4} 2f(x) + 1/2 dx \]

OpenStudy (kc_kennylau):

Split it into two integrals

OpenStudy (anonymous):

hmm, ok. i will try.

OpenStudy (anonymous):

\[\int\limits_{-1}^{4} 2f(x) dx \] + \[\int\limits_{-1}^{4} 1/2 dx \]

OpenStudy (anonymous):

is it like this?

OpenStudy (kc_kennylau):

Yes

OpenStudy (anonymous):

the f(x) dx, I must put 3 ?

OpenStudy (kc_kennylau):

It's 2f(x)dx

OpenStudy (kc_kennylau):

So it's 6

OpenStudy (anonymous):

i dont get the answer,

OpenStudy (kc_kennylau):

This is not the answer

OpenStudy (kc_kennylau):

But \(\Large\displaystyle\int_{-1}^42f(x)dx=2\int_{-1}^4f(x)dx=2\times3=6\)

OpenStudy (anonymous):

and then it become 6x ? and other one become 1/2 x

OpenStudy (anonymous):

@ganeshie8 can you help me out here ?

ganeshie8 (ganeshie8):

\(\large \int\limits_{-1}^{4} 2f(x) + \frac{1}{2} dx\) \(\large \left(2\int\limits_{-1}^{4} f(x) dx\right) + \left(\int\limits_{-1}^{4} \frac{1}{2} dx\right)\) \(\large \left(2\times 3\right) + \left(\frac{1}{2} \int\limits_{-1}^{4} 1dx\right)\) \(\large \left(6\right) + \left(\frac{1}{2} [4--1]\right)\) \(\large \left(6\right) + \left(\frac{5}{2} \right)\) \(\large \frac{17}{2} \)

ganeshie8 (ganeshie8):

let me knw if somthng doesnt make sense..

OpenStudy (anonymous):

let me try first..

ganeshie8 (ganeshie8):

okiee

OpenStudy (anonymous):

ok, completely understand... :D thank you again. I glad you are around when I need ya.

ganeshie8 (ganeshie8):

good to hear :) u wlc !!